OpenStudy (anonymous):

a and b are the lengths of the legs of a right triangle, and c is the hypotenuse. Find the missing length if a=15 and c=39... How to do this?

OpenStudy (anonymous):

Using the Pythagorean Theorem, of course :) The hypotenuse c of any right triangle can be calculated using a pretty nifty formula \[\huge a^2 + b^2 = c^2\] Where a and b are the legs...

OpenStudy (anonymous):

So would it be 225 + b2 = 1521 after plugging it in and the work?

OpenStudy (anonymous):

Yes, now just work out b^2 and consequently, work out b :)

OpenStudy (anonymous):

Ok but this is where im a little confused, how would i "work out b"? lol I dont plug anything into that right? Lol

OpenStudy (anonymous):

You don't. think of it as a variable that you're solving for.

OpenStudy (anonymous):

Ok thanks for the help!

OpenStudy (anonymous):

^.^